# MSRS: Adaptive Multi-Subspace Representation Steering for Attribute Alignment in Large Language Models

### v3

@@ 4.3 Optimization Objective
We further encourage the learned representation $R$ to align with the structured basis $S_{\text{align}}$, as defined in equation 2. The alignment loss is defined as $\mathcal{L}_{\text{align}}=1-\frac{\langle R,S_{\text{align}}\rangle}{\|R\|_{2}\|S_{\text{align}}\|_{2}}$, which encourages $R$ to lie in the subspace spanned by both shared and attribute-specific directions, promoting more controllable and semantically meaningful representations during training. The overall optimization objective is as follows:
@@
where $\lambda_{1},\lambda_{2}>0$ are hyperparameters balancing the terms. This optimization guarantees both attribute-wise subspace alignment and inter-attribute separation, ultimately yielding an effective steering space capable of precise and disentangled multi-attribute control. By integrating different subspaces with a weighting network, we enable adaptive subspace combination, alleviating trade-offs and optimizing performance across diverse attributes’ steering. See Appendix B for the detailed algorithm.
4.4 Dynamic Intervention Position Selection
Previous steering approaches typically apply interventions at the same token positions (often at the suffix) for different attributes, leading to suboptimal positioning for attribute steering, which hinders optimal results wang2024helpsteer. To overcome this limitation, we propose a dynamic selection method that identifies the most relevant token position $p_{i}$ for each attribute $i$ by projecting token representations from the last few tokens onto attribute-specific subspaces in $\mathbb{R}^{d}$. This enhances steering effectiveness for attribute $i$ by targeting interventions at the most influential tokens.
Consider an input sequence with token representations $h_{1},h_{2},\dots,h_{T}$, where $T$ denotes the sequence length. For each attribute $i$, we project the token representations onto its corresponding subspace $R_{i}$. The projection of a token representation $h_{t}$ onto this subspace is computed as $\text{proj}_{R_{i}}(h_{t})=R_{i}^{\top}R_{i}h_{t}$. We then define the relevance score $s_{i,t}$ of token $t$ with respect to attribute $i$ as the $L_{2}$-norm of this projection: $s_{i,t}=\|\text{proj}_{R_{i}}(h_{t})\|_{2}$. The intervention position $p_{i}$ for attribute $i$ is dynamically selected as $p_{i}=\arg\max_{t\in\{1,\dots,T\}}s_{i,t}$, ensuring that the token with the strongest alignment to the attribute-specific subspace is chosen. The steering function $\Phi_{l,p}(h;R,W,b,m)$ for attribute $i$ is applied at position $p_{i}$, allowing different attributes to be steered at different tokens, reducing inter-attribute interference and enhancing control precision.
5 Experiments
5.1 Experimental Settings
Datasets and Metrics. To evaluate the effectiveness of our proposed method, we conduct experiments on three pairs of datasets, each designed to assess trade-offs between specific attributes in language model steering. More details are provided in Appendix 1. 1) TruthfulQA & BBQ: To evaluate truthfulness and bias, we use TruthfulQA with MC1, MC2, BLEU, and BLEURT scores (Lin et al., 2022), and BBQ with accuracy as the metric (Parrish et al., 2022). 2) Alpaca & Refusal: We evaluate instruction following via win rate on Alpaca (Taori et al., 2023; Li et al., 2023b) (vs. test-davinci-003), and refusal via Sorry-Bench scores judged by Mistral-7B-Instruct-v0.2 (Xie et al., 2025), which assesses the rejection of malicious instructions. 3) HelpSteer: We assess helpfulness, coherence, and verbosity by leveraging GPT-3.5-Turbo, following the setting of (Nguyen et al., 2025a), to rate model outputs on a 0–4 scale (Wang et al., 2023), which evaluates the quality of the generated content.
Additionally, we test the utility on several standard benchmarks, including Hellaswag (Zellers et al., 2019), RACE (Lai et al., 2017), MMLU (Hendrycks et al., 2020), OpenBookQA (Mihaylov et al., 2018), and GLUE (Wang et al., 2018), all using accuracy as the evaluation metric. We aim to assess whether MSRS preserves the model’s general abilities after fine-tuning, ensuring that task-specific steering does not degrade overall performance.
Table 1: Evaluation results on TruthfulQA, BBQ, Alpaca, Refusal, and HelpSteer, with the best result in bold and the second-best underlined. The reported values are mean performances.
@@ 5.1 Experimental Settings
[…]
Models and Baselines. We evaluate MSRS on 4 models: Llama2-7B (Touvron et al., 2023), Llama3-8B-Instruct (Grattafiori et al., 2024), Qwen2-7B-Instruct (Team, 2024) , Mistral-7B-v0.3 (Jiang et al., 2023). And we compare MSRS with 6 baselines, grouped into 3 categories: 1) In-context Learning (Brown et al., 2020): Utilizes prompts to steer attributes without altering model parameters. 2) Fine-tuning Methods: MTL-LoRA (Yang et al., 2025c), which employs low-rank adaptation for multi-task learning, enabling efficient attribute-specific fine-tuning. ReFT (Wu et al., 2024b), which adjusts model representations by fine-tuning the representation to align with target attributes. 3) Steering Methods: ITI (Li et al., 2023a) applies inference-time interventions to modify activations and guide model outputs. CAA (Rimsky et al., 2024) steers behavior by injecting contrastive activation vectors derived from positive and negative examples. MAT-STEER (Nguyen et al., 2025b) implements multi-attribute steering with orthogonal constraints to minimize interference between attributes.
Experimental Setup. All experiments were conducted on NVIDIA V100 GPUs using the Adam optimizer with a learning rate of $5\times 10^{-3}$. The total subspace rank $R$ was 8, with dual regularization coefficients $\lambda_{1}=0.3$, $\lambda_{2}=0.5$. Steering was applied to the 15th transformer layer, selected as the most effective on the validation set. For each configuration, we report the average and standard deviation over 3 runs with random seeds {42, 43, 44}. In Appendix F, we provide a description of the hyperparameter selection method and sensitivity analysis to assess the robustness of MSRS to these hyperparameters, along with the computational cost for different configurations.
@@ 5.2 Main results
MSRS demonstrates strong performance in open-ended generation tasks. We assess MSRS on Alpaca, Refusal, and HelpSteer datasets to evaluate trade-offs between instruction following, refusal, and output quality attributes. As shown in Table 1, MSRS excels in balancing instruction-following and refusal. On Alpaca, it achieves a win rate of 0.36 against test-davinci-003 on Llama3-8B-Instruct, outperforming ReFT (0.30), and on Refusal, it scores 0.529 on Sorry-Bench, surpassing CAA’s 0.493. Baselines like ITI struggle with refusal (0.280), sacrificing this attribute for others. For HelpSteer, MSRS consistently improves Helpfulness, Coherence, and Verbosity on Llama3-8B-Instruct, while methods like MAT-STEER improve Helpfulness but reduce Coherence and Verbosity. MTL-LoRA achieves the best Verbosity, but at the cost of Helpfulness. MSRS outperforms existing baselines in harmonizing instruction-following with refusal and delivers high-quality generations across all HelpSteer dimensions, demonstrating its ability to integrate diverse attributes and enhance open-ended generation for specific attributes. Even when MSRS is not the absolute best (e.g., Coherence on Mistral-7b-v0.3), it achieves near-best results with a negligible gap and retains overall superiority across all attributes.
@@
Method HellaSwag RACE MMLU OpenbookQA GLUE
Llama3-8b-inst. 0.801 0.671 0.655 0.556 0.726
ReFT 0.821 0.677 0.651 0.559 0.757
ITI 0.746 0.589 0.546 0.507 0.742
MTL 0.782 0.661 0.567 0.562 0.697
CAA 0.833 0.671 0.648 0.557 0.738
[…]
Qwen2-7b-inst. 0.831 0.625 0.695 0.606 0.825
ReFT 0.822 0.644 0.698 0.613 0.770
MTL 0.782 0.641 0.687 0.562 0.697
CAA 0.837 0.633 0.698 0.609 0.830
[…]
Mistral-7b-v0.3 0.862 0.678 0.618 0.602 0.681
ReFT 0.872 0.679 0.603 0.608 0.655
MTL 0.869 0.644 0.530 0.574 0.682
CAA 0.869 0.667 0.619 0.611 0.693
[…]
MSRS maintains strong general capabilities on standard NLP benchmarks. To verify that MSRS does not compromise the model’s overall natural language processing abilities, we evaluate its performance on several widely used benchmarks, as shown in Table 2. More results on the MMLU (Table 6) and GLUE benchmark tasks (Table 7) are shown in Appendix  B.1. These tasks assess a model’s general reasoning, knowledge, and language understanding, offering a comprehensive measure of its robustness beyond attribute-specific steering. From the results, baseline methods often compromise general capabilities by overfitting to specific attributes. For example, ITI performs well on truthfulness but struggles on general benchmarks like HellaSwag and MMLU on LLaMA3, showing weakened commonsense and knowledge reasoning. In contrast, MSRS consistently matches or exceeds baseline performance across these benchmarks. For GLUE, MSRS scores an average of 0.775, surpassing ITI and ReFT. Figure 3 compares different tasks in GLUE, where MSRS shows substantial improvements on SST-2 and STS-B, reflecting better sentiment classification and semantic similarity. This robust performance is driven by MSRS’s shared subspace mechanism, which captures common steering directions across tasks and attributes, enhancing generalization and excelling in both general NLP benchmarks and targeted steering objectives.
5.3 Ablation Study
The shared subspace facilitates the integration of multi-attribute features, resulting in enhanced overall performance. Since the effectiveness of MSRS depends on the shared subspace, we investigate the impact of the rank ($k$ of the shared subspace) on performance. As shown in Figure 3, we report the relationship between the performance (Avg. Accuracy) of the Truthfulness and Bias attributes and the proportion of the shared subspace within the total space (shared rank ratio).
Figure 2: Comparison of model performance across GLUE.
@@ 5.3 Ablation Study
Models with a shared subspace in $R$ outperform those without shared subspace (w/o Shared Subspace, only attribute-specific space) or with no subspace partition. The optimal shared subspace ratio varies with $R$: for $R=8$, it is around 30%, and for $R=12$, it is about 20%. As the shared subspace ratio increases, the attribute-specific space decreases, causing performance to decline. At 100%, the model degenerates to the no subspace partition case. Additionally, we display the explained standard deviation (blue bars, i.e., the ratio of preserved singular values $\sigma$ to the total sum of singular values $\Sigma$). The optimal shared subspace ratio is between 20% and 40%, where the explained standard deviation reaches approximately 60%. Accordingly, we choose the fewest singular vectors that capture at least 60% of the total energy in $\Sigma_{c}$. See additional analysis in Appendix G.
Table 3: Comparison of steering subspace training strategies across datasets.
Method TruthfulQA BBQ Alpaca Refusal HelpSteer
MC1 MC2 Acc Win Sorry Help. Coh. Ver.
Llama3-8b-inst.
Same Space 29.58 49.51 0.637 0.30 0.451 3.87 3.89 2.38
@@
Qwen2-7b-inst.
Same Space 29.83 48.69 0.637 0.434 0.422 3.63 3.78 2.38
[…]
Mistral-7b-v0.3
Same Space 30.32 49.69 0.615 0.33 0.669 3.82 3.85 2.33
[…]
Table 4: Comparison of Last token vs. Important token intervention.
Method TruthfulQA BBQ Alpaca Refusal HelpSteer
MC1 MC2 Acc Win Sorry Help. Coh. Ver.
LLaMA2-7B
Last Token 26.41 42.88 0.631 0.12 0.579 2.70 2.68 2.73
Important Token 29.10 48.60 0.644 0.13 0.583 3.12 3.06 2.47
LLaMA3-8B-Inst.
Last Token 33.50 52.74 0.648 0.36 0.529 3.88 3.96 2.24
Important Token 33.71 56.32 0.655 0.32 0.511 3.85 3.95 1.99
Qwen2-7B-Inst.
Last Token 34.72 53.27 0.6421 0.45 0.446 3.70 3.82 2.17
Important Token 36.12 55.63 0.6572 0.42 0.448 3.69 3.83 2.06
Evaluating the effectiveness of adaptive subspace selecting mechanism. We conduct ablation studies comparing three strategies for training steering subspaces: (1) Same Space, where all attributes are trained in a single subspace without isolation; (2) $\text{MSRS}_{\text{Attribute}}$, the basis matrix $R\in\mathbb{R}^{r\times d}$ is partitioned into $n+1$ blocks: $R=\left[B_{\text{shared}}\;\middle\|\;B_{1}\;\middle\|\;\dots\;\middle\|\;B_{n}\right]$, where $B_{\text{shared}}$ is a shared subspace and each $B_{i}$ corresponds to a specific attribute. The mask network $m(h)$ generates soft weights that are applied to each block individually, allowing the model to adaptively activate relevant attribute-specific subspaces; and (3) $\text{MSRS}_{\text{Rank}}$ treats the basis matrix $R\in\mathbb{R}^{r\times d}$ as a flat set of $r$ independent basis vectors, without any explicit block structure. The mask network $m(h)\in\mathbb{R}^{r}$ assigns a soft weight to each row of $R$, enabling control at the level of individual basis directions. In our implementation, we adopt the $\text{MSRS}_{\text{Attribute}}$ configuration as the default setup for MSRS. This choice offers a good balance between interpretability and control, allowing the model to modulate behavior based on attribute-specific subspaces. Results are reported in Table 4 across multiple datasets and models.
Training all attributes in the same space leads to suboptimal outcomes due to interference between conflicting objectives. For example, Same Space performs worse than both MSRS variants across nearly all metrics. MSRS mitigates this by decoupling subspaces and adaptively selecting them via $m(h)$. The $\text{MSRS}_{\text{Attribute}}$ configuration, which groups low-rank dimensions into attribute-aligned subspaces, strikes a balance between parameter sharing and specialization, consistently improving both truthfulness and bias mitigation across models. The $\text{MSRS}_{\text{Rank}}$ variant offers finer-grained control, improving performance on LLaMA3-8B-Instruct and suggesting better truthfulness. However, it lags behind $\text{MSRS}_{\text{Attribute}}$ on Qwen2-7B-Instruct, possibly due to differing optimal subspace granularity across models, with some benefiting from coarser grouping. Additional results are in Appendix C (Table 8).
Validating the effectiveness of proposed dynamic intervention position selection mechanism. We compare two settings: (1) the Last token, where intervention is applied to the last token in the sequence, and (2) Important token, which applies intervention at the dynamically selected Important token from the last few tokens. As shown in Table 4, dynamic token selection consistently outperforms fixed-position steering across all models. For example, on LLaMA2-7B, MC1 improves from 26.41 to 29.10 and BBQ accuracy from 0.631 to 0.644. Similar gains are observed on LLaMA3-8B-Instruct and Qwen2-7B-Instruct. These results validate that selecting the most semantically relevant token for intervention mitigates inter-attribute interference and enables more effective attribute control. Additional results are shown in Appendix  D (Table 9).
Explore the relationship between the $R$-similarity of selected steering token and model performance. We designed an experiment where the intervention was applied to different single token at the suffix position, as illustrated in Figure 4. The results indicate that steering at the token with higher similarity to $R$ generally leads to better performance. Moreover, this intervention consistently
@@
Figure 4: Token–$R$ similarity vs. performance under different token intervention. Higher similarity correlates with better performance.
outperforms the no-intervention baseline across all tokens, demonstrating its generalization ability across different tokens. We attribute this consistent improvement to using the dominant directions of average activations from the last few tokens to guide training, enabling the learned $R$ to generalize to some extent to different tokens. By dynamically selecting tokens for intervention during inference based on their similarity to $R$, different attributes can be applied to those tokens most closely related to the desired attribute steering, yielding more precise control over attributes.
Figure 5: Performance of interventions at different transformer layers. Mid-layer intervention consistently outperforms others.
Determination of the optimal layer for steering interventions. We evaluate model performance by injecting steering vectors at different transformer layers, applying interventions only at each candidate layer, as shown in Figure 5. Lower layers show weak steering likely due to insufficient semantic abstraction, while mid-to-upper layers generally yield better results, with layer 15 achieving the best performance. Deeper layers, however, tend to overfit, leading to performance degradation. These findings align with prior work (skean2025layer), which suggests that mid-depth layers in transformer architectures play a key role in controlling downstream tasks. Detailed results for different layers are in Appendix E (Table 10). To select the optimal intervention layer for multi-attribute subspace training, we conduct a grid search over held-out validation splits to identify the layer that best balances attribute trade-offs. This strategy guides our selection of the intervention layer for all subsequent experiments, ensuring interventions are applied where most effective, thereby maximizing the utility of the learned multi-subspace representations.
Table 5: Ablation study on multi-subspace alignment strategies.
Method MC1 MC2 BLEU BLEURT BBQ (Acc)
LLaMA2 18.58 35.25 38.37 52.65 0.604
+ReFT 21.03 36.93 44.74 55.99 0.626
+CS-ReFT 22.85 36.12 44.56 65.47 0.634
+ReFT + SVD 22.58 38.97 46.45 58.92 0.633
[…]
Evaluating the benefit of aligning learned steering subspaces with SVD-derived priors. We conduct an ablation study on LLaMA2-7B using four configurations: (1) standard ReFT, which learns a single steering subspace without partitioning; (2) CS-ReFT (Zhou, 2025), which partitions the total rank $R$ into equal-sized subspaces and learns a weighted matrix to combine them; (3) Naive SVD-based subspace concatenation, which applies SVD to extract low-rank basis vectors for each attribute and concatenates them with the shared subspace basis to initialize the projection matrix $R$ in ReFT; (4) Subspace alignment strategy, which aligns the learned subspace $R$ with SVD priors $m_{\text{prior}}$ during training using a mask regularization loss $\mathcal{L}_{\text{reg}}$ and a directional alignment loss $\mathcal{L}_{\text{align}}$, enabling adaptive, interpretable, and task-aware control. The results are shown in Table 5. The subspace alignment strategy (+ReFT+Align) outperforms both baselines across all metrics, surpassing ReFT and SVD-only approaches. Compared to CS-ReFT, our method allocates subspaces more effectively by aligning them with SVD priors, resulting in improved performance. These results highlight the importance of incorporating structural priors during training, with SVD priors serving as an effective foundation for guiding subspace learning.

### v4

@@ 4.3 Optimization Objective
We further encourage the learned representation $R$, which is initialized to match the dimensions of $S_{\text{align}}$, to align with the structured basis $S_{\text{align}}$, as defined in (3). The alignment loss is defined as $\mathcal{L}_{\text{align}}=1-\frac{\langle R,S_{\text{align}}\rangle_{F}}{\|R\|_{F}\|S_{\text{align}}\|_{F}}$, where $\langle A,B\rangle_{F}=\text{Tr}(A^{\top}B)=\sum_{i,j}A_{ij}B_{ij}$ denotes the Frobenius inner product, and $\|\cdot\|_{F}$ represents the Frobenius norm. This formulation encourages $R$ to lie in the subspace spanned by both shared and attribute-specific directions, promoting more controllable and semantically meaningful representations during training.
The overall optimization objective is as follows:
@@
where $\lambda_{1},\lambda_{2}>0$ are hyperparameters balancing the terms. This optimization guarantees both attribute-wise subspace alignment and inter-attribute separation, ultimately yielding an effective steering space capable of precise and disentangled multi-attribute control. By integrating different subspaces with a weighting network, we enable adaptive subspace combination, alleviating trade-offs and optimizing performance across diverse attributes’ steering.
Table 1: Evaluation results on TruthfulQA, BBQ, Alpaca, Refusal, and HelpSteer.
@@
MSRS 31.32±0.11 52.62±0.13 50.61±0.08 71.39±0.17 0.644±0.007 0.38±0.006 0.693±0.013 3.82±0.02 3.93±0.01 2.21±0.02
5 Experiments
5.1 Experimental Settings
Datasets and Metrics. We evaluate MSRS on three dataset pairings designed to assess multi-attribute steering trade-offs (details in Appendix A): 1) TruthfulQA & BBQ: We measure truthfulness via TruthfulQA metrics (MC1/2, BLEU, BLEURT) (Lin et al., 2022) and social bias via BBQ accuracy (Parrish et al., 2022). 2) Alpaca & Refusal: We evaluate instruction following using Alpaca win-rates (vs. text-davinci-003) (Taori et al., 2023; Li et al., 2023b) and safety refusal via Sorry-Bench scores (judged by Mistral-7B-v0.2) (Xie et al., 2025). 3) HelpSteer: We assess helpfulness, coherence, and verbosity on a 0–4 scale using a GPT-4-o judge (Nguyen et al., 2025; Wang et al., 2023). To ensure steering does not degrade general capabilities, we also report General Utility accuracy across standard benchmarks: Hellaswag (Zellers et al., 2019), RACE (Lai et al., 2017), MMLU (Hendrycks et al., 2020), OpenBookQA (Mihaylov et al., 2018), and GLUE (Wang et al., 2018).
Models and Baselines. We evaluate MSRS on 4 models: Llama2-7B (Touvron et al., 2023), Llama3-8B-Instruct (Grattafiori et al., 2024), Qwen2-7B-Instruct (Team, 2024) , Mistral-7B-v0.3 (Jiang et al., 2023). And we compare MSRS with 6 baselines, grouped into 3 categories: 1) In-context Learning (Brown et al., 2020): Utilizes prompts to steer attributes without altering model parameters. 2) Fine-tuning Methods: ReFT (Wu et al., 2024), which adjusts model representations by fine-tuning the representation to align with target attributes. CS-ReFT (Zhou, 2025), which employs a lightweight router to compose task-specific orthonormal subspaces for interference-free adaptation. 3) Steering Methods: ITI (Li et al., 2023a) applies inference-time interventions to modify activations and guide model outputs. CAA (Rimsky et al., 2024) steers behavior by injecting contrastive activation vectors derived from positive and negative examples. MAT-STEER (Nguyen et al., 2025) implements multi-attribute steering with orthogonal constraints to minimize interference between attributes.
Experimental Setup. All experiments were conducted on NVIDIA V100 GPUs. We employed the Adam optimizer with a learning rate of $9\times 10^{-4}$ and a batch size of 4. The total subspace rank $R$ is set to match the rank of $S_{\text{align}}$ , with dual regularization coefficients $\lambda_{1}=0.3$, $\lambda_{2}=0.5$ (robustness verified in Appendix F). Following the ReFT setting, we fixed the intervention to the last 5 tokens. For each configuration, we report the average and standard deviation over 3 runs with random seeds {42, 43, 44}.
@@ 5.2 Main results
MSRS demonstrates strong performance in open-ended generation tasks. We assess MSRS on Alpaca, Refusal, and HelpSteer datasets to evaluate trade-offs between instruction following, refusal, and output quality attributes. As shown in Table 1, MSRS demonstrates superior capability in balancing instruction-following and refusal. On Llama3-8B-Instruct, it effectively surpasses strong baselines like ReFT on Alpaca and CAA on Sorry-Bench, whereas methods like ITI suffer from significant degradation in refusal performance. Regarding HelpSteer, MSRS consistently enhances Helpfulness, Coherence, and Verbosity. In contrast, other baselines exhibit clear trade-offs, prioritizing specific metrics (e.g., Verbosity) at the expense of others. Overall, MSRS achieves comprehensive superiority; even in rare cases where it does not rank first, it maintains comparable performance with a negligible gap, avoiding the severe performance drops seen in other methods.
Figure 3: Comparison of model performance across GLUE.
@@
Model / Method HellaSwag RACE MMLU OBQA GLUE
Llama-3-8B-Instruct
     Vanilla 0.801 0.671 0.655 0.556 0.726
     + MSRS (Ours) 0.839 0.683 0.657 0.568 0.775
Qwen2-7B-Instruct
     Vanilla 0.831 0.625 0.695 0.606 0.825
     + MSRS (Ours) 0.835 0.648 0.702 0.616 0.832
Mistral-7B-v0.3
     Vanilla 0.862 0.678 0.618 0.602 0.681
     + MSRS (Ours) 0.874 0.681 0.613 0.622 0.707
MSRS maintains strong general capabilities on standard NLP benchmarks. We evaluate MSRS on standard benchmarks (Table 2 and Appendix A.1) to assess its robustness. MSRS preserves the base model’s utility, consistently achieving comparable or superior performance. Specifically, on the GLUE benchmark, MSRS attains an average score of 0.775, outperforming competing methods (see details in Appendix  A.2). Figure 3 further highlights improvements in sentiment and semantic tasks. This superior generalization is driven by our shared subspace mechanism, which learns common steering directions applicable across tasks, thereby avoiding the negative trade-offs distinct in prior works.
@@
5.3 Core Multi-Subspace Design
The shared subspace facilitates the integration of multi-attribute features, resulting in enhanced overall performance. Since the effectiveness of MSRS depends on the shared subspace, we investigate the impact of the rank ($k$ of the shared subspace) on performance, by exploring the allocation of the shared subspace under the condition of the same total subspace rank $R$. As shown in Figure 4, we report the relationship between the performance (Avg. Accuracy) of the attributes and the proportion of the shared subspace within the total space (shared rank ratio).
Models with a shared subspace in $R$ outperform those without shared subspace (w/o Shared Subspace, only attribute-specific space) or with no subspace partition. The optimal shared subspace ratio varies with $R$: for $R=8$, it is around 30%, and for $R=12$, it is about 20%. As the shared subspace ratio increases, the attribute-specific space decreases, causing performance to decline. At 100%, the model degenerates to the no subspace partition case. Additionally, we display the explained standard deviation (blue bars, i.e., the ratio of preserved singular values $\sigma$ to the total sum of singular values $\Sigma$). The optimal shared subspace ratio is between 20% and 40%, where the explained standard deviation more than 60%. Accordingly, we choose the fewest singular vectors that capture at least 60% of the total energy in $\Sigma_{c}$. See sensitivity analysis in Appendix G.
Table 3: Comparison of steering subspace training strategies.
TruthfulQA BBQ Alpaca Refusal HelpSteer
Method Avg. Acc Win Sorry Avg.
Same Space 39.55 0.637 0.30 0.451 3.88
@@ 5.3 Core Multi-Subspace Design
Mechanism Analysis. We compare three strategies for training steering subspaces: (1) Same Space: All attributes are trained within a single global subspace without isolation. (2) $\text{MSRS}_{\text{Attribute}}$ (Default): The basis matrix $R$ is partitioned into semantic blocks $R=[B_{\text{shared}}\,|\,B_{1}\,|\,\dots\,|\,B_{n}]$, where $m(h)$ modulates each block to adaptively activate specific attribute subspaces. (3) $\text{MSRS}_{\text{Rank}}$: The basis $R$ is treated as $r$ independent vectors, with $m(h)\in\mathbb{R}^{r}$ weighting each rank individually for fine-grained control. We adopt $\text{MSRS}_{\text{Attribute}}$ as the default implementation, as it offers the optimal balance between interpretability and effective attribute-level steering. As shown in Table 3, the Same Space baseline underperforms due to interference from conflicting objectives within a shared optimization space. MSRS mitigates this through subspace decoupling and adaptive modulation $m(h)$. Specifically, $\text{MSRS}_{\text{Attribute}}$ uses coarse-grained partitioning with dedicated attribute blocks, achieving consistent efficiency and gains in instruction-following tasks (e.g., Alpaca, Refusal). In contrast, $\text{MSRS}_{\text{Rank}}$ provides fine-grained control by modulating individual basis directions, yielding superior performance in tasks requiring precise feature suppression, such as Truthfulness and BBQ. (See Appendix C, Table 10 for full results).
5.4 Enhanced Steering Mechanisms
@@
Figure 5: Token–$R$ similarity vs. performance under different token steering. Higher similarity correlates with better performance.
Figure 6: Comparison of Last token vs. Important token steering.
Figure 7: Performance of interventions at different layers.
Table 4: Inference-time comparison of different steering methods.
Method Latency (ms/tok) $\downarrow$ Throughput (tok/s) $\uparrow$ VRAM (GB) $\downarrow$
Baseline 102.44 9.76 16.10
ReFT 115.72 8.64 16.26
CAA 110.59 9.04 16.12
MSRS (Ours) 120.35 8.38 16.29
Projection similarity serves as a proxy for steerability. Our core insight is that steering efficacy varies across token positions; tokens aligning closely with the attribute-specific subspace ($R$) are semantically relevant, serving as optimal anchor points for intervention. To validate this, we examined the correlation between $R$-similarity and model performance by applying interventions at distinct single-token suffix positions (Figure 5). The results empirically confirm our hypothesis: steering at tokens with higher similarity is more likely to yield superior performance. Moreover, the intervention outperforms the no-intervention baseline across all positions, demonstrating robust generalization.
To quantify this capability, we conducted a comparative analysis between two intervention strategies: (1) Last Token, where the steering vector is applied strictly to the final token of the sequence; and (2) Important Token, which dynamically selects the intervention point from the recent context based on the highest projection similarity. As shown in Figure 6, the dynamic Important Token strategy demonstrates competitive performance gains over the fixed Last Token baseline. This indicates that the attribute-specific subspace derived by MSRS possesses the intrinsic capability to identify optimal anchor points for attribute intervention.
Optimal Layer Selection. We evaluate steering efficacy across different layers of LLaMA3 (Figure 7). Consistent with prior findings (Skean et al., 2025), mid-to-upper layers yield superior control, whereas lower layers lack semantic abstraction and deeper layers are prone to overfitting. Consequently, based on a grid search balancing attribute trade-offs on validation splits, we select layer 15 for LLaMA3, layer 9 for Qwen2, and layer 21 for Mistral for all subsequent experiments. Detailed results are provided in Appendix E.
Computational Efficiency. We evaluated the computational overhead of various methods using Llama3-8B-Instruct on the Alpaca dataset (sequence length 512, batch size 1). As shown in Table 4, MSRS introduces modest latency and VRAM overhead compared to the base model, with performance metrics remaining comparable to ReFT and only marginally higher than CAA. Appendix B further confirms the minimal overhead of the one-time offline initialization phase. Given that MSRS consistently delivers the superior
Figure 8: Impact of Data Contamination on Subspace Semantics.
performance across all the attributes, this minor overhead represents a favorable trade-off between operational efficiency and multi-attribute steering capability.
5.5 Sensitivity to Data Coupling
To assess the robustness of MSRS against data contamination arising from attribute co-occurrence (where target samples inextricably contain distractor features), we investigate how data-level coupling transmits to representation-level entanglement. We simulated varying degrees of contamination by blending a pure target dataset (Helpfulness) with a distractor dataset (Coherence) at a mixing ratio $\alpha$. We trained steering matrices on these mixed sets and evaluated their semantic alignment using a Relative Entanglement Score (RES), which measures the subspace’s projection ratio between the distractor and the target oracle (detailed setup and metric definitions are provided in Appendix H). As shown in Figure 8, MSRS exhibits a distinct non-linear robustness profile. In the Robustness Regime ($\alpha\leq 0.55$), the method demonstrates remarkable resilience: even when distractor data constitutes the majority, the learned subspace maintains low entanglement ($\text{RES}<0.62$) and stable steering performance ($12.5\%$). A critical Breakdown Point is observed at $\alpha\approx 0.61$, where the RES surges and performance drops sharply, indicating a ”tipping point” where distractor gradients begin to dominate. Notably, even under severe contamination ($\alpha=0.80$), MSRS retains positive steering efficacy ($1.5\%$), confirming its intrinsic capability to filter data noise and extract discriminative features, validating the robustness of our subspace-based intervention.
Figure 9: Normalized Relative Percentage Change ($\Delta\%$) in Different Attributes Compared to Baseline (n = 0).
Figure 10: Interference, Performance, and Rank(k) vs. Number of Attributes(n).
5.6 Scalability and Interference Patterns with Growing Attribute Count
Investigate the relationship between the number of attributes $n$ and performance in multi-attributes scenario. Figure 9 visualizes the relative performance change ($\Delta\%$) for each attribute $a_{i}$ ($i\in\{1,\dots,10\}$) across $n\in\{0,\dots,10\}$ attributes (attributes details are shown in Appendix A), relative to the baseline ($n=0$). For each attribute $a_{i}$ and number of attributes $n=j$, we construct $M$ random subsets $S_{i,j}^{m}$ ($m\in\{1,\dots,M\}$), each defined as $S_{i,j}^{m}=\{a_{i}\}\cup T_{i,j}^{m}$, where $T_{i,j}^{m}\subset\mathcal{A}\setminus\{a_{i}\}$ is a random subset of j1j-1 attributes sampled from the 9 attributes excluding $a_{i}$, with $|\mathcal{A}|=10$; in our setup, $M=5$. The performance of $a_{i}$ in $j$-joint attributes is measured after joint training, and averaged as $\text{Perf}_{a_{i}}(n=j)=\frac{1}{M}\sum_{m=1}^{M}\text{Perf}_{a_{i}}(S_{i,j}^{m})$. The baseline is $\text{Perf}_{a_{i}}(n=0)$, and the relative performance change is given by $\Delta_{a_{i}}(n=j)=\frac{\text{Perf}_{a_{i}}(n=j)-\text{Perf}_{a_{i}}(n=0)}{\text{Perf}_{a_{i}}(n=0)}$. The $\Delta_{a_{i}}(n)$ is normalized to [0, 1] as $\text{Norm}_{\Delta_{a_{i}}}(n)=\frac{\Delta_{a_{i}}(n)}{\max_{n}\Delta_{a_{i}}(n)-\min_{n}\Delta_{a_{i}}(n)}$. The heatmap shows $\text{Norm}_{\Delta_{a_{i}}}(n)$ as color intensity (red: negative, blue: positive) with $\Delta_{a_{i}}(n)$ annotated (See Table 6 for the performance results). As shown in Figure 9, for n=24n=2-4, most attributes show strong improvements ($\Delta\%>0$, deep blue), indicating effective attributes joint. At n=58n=5-8, gains weaken (lighter blue), reflecting uneven benefits. At n=910n=9-10, some attributes drop to or below baseline ($\Delta\%\leq 0$, red/neutral), signaling performance degradation. We attribute these trends to interference in the shared subspace driven by increasing attribute counts. At low $n$, attributes align with minimal interference. As $n$ grows, dominant attributes (e.g., Refusal, Style) with positive $\Delta\%$ bias the shared subspace, their steering vectors overpowering weaker attributes, which exhibit sharply reduced gains. At n=910n=9-10, intensified interference causes performance declines across more attributes.
Performance, interference, and rank vary with number of attributes $n$ in multi-attribute steering. Figure 10 illustrates the relationships between average performance (green), interference (blue), and rank $k$ (red) with $n$. Performance is computed as the mean over all attributes. Interference quantifies disorder in the shared subspace, manifested as consistent performance declines across attributes. It is defined as the mean normalized decline from peak performance: for each attribute $a_{j}$, the decline is given by $\text{Decline}_{a_{j}}(n)=\frac{\text{PeakPerf}_{a_{j}}-\text{Perf}_{a_{j}}(n)}{\text{PeakPerf}_{a_{j}}}$, with $\text{PeakPerf}_{a_{j}}=\max_{n}\text{Perf}_{a_{j}}(n)$ as the highest performance and $\text{Perf}_{a_{j}}(n)$ as the performance at $n$. The average interference is computed as $\text{AvgInterference}(n)=\frac{1}{m}\sum_{j=1}^{m}\text{Decline}_{a_{j}}(n)$, where $m=10$ is the number of attributes.
The trends approximate a quadratic fit for performance ($R^{2}=0.9520$), a linear fit for rank $k$ ($R^{2}=0.9427$), and a cubic fit for interference ($R^{2}=0.9903$) (see Appendix I for fit parameters). The green performance curve shows sustained improvements over baseline for $n<9$, peaking at $n=2$ or 3 with subsequent gradual reductions in gains. The blue curve indicates accelerating interference with $n$, while optimal performance corresponds to rank $k=8$ to 12, consistent with Table 14. These patterns suggest that at low $n$, attributes align synergistically, enabling superior multi-attribute steering performance and demonstrating attribute scalability for $n<9$, corresponding $\text{AvgInterference}<0.6$.